# Read Between the Lines: A Benchmark for Uncovering Political Bias in Bangla News Articles

### v1

@@ Abstract
Detecting media bias is crucial, specifically in the South Asian region. Despite this, annotated datasets and computational studies for Bangla political bias research remain scarce. Crucially because, political stance detection in Bangla news requires understanding of linguistic cues, cultural context, subtle biases, rhetorical strategies, code-switching, implicit sentiment, and socio-political background. To address this, we introduce BanglaBias, the first benchmark dataset of 200 politically significant and highly debated Bangla news articles, labeled for government-leaning, government-critique, and neutral stances, alongside diagnostic analyses for evaluating large language models (LLMs). Our comprehensive evaluation of 28 proprietary and open-source LLMs shows strong performance in detecting government-critique content (F1 up to 0.83) but substantial difficulty with neutral articles (F1 as low as 0.00). Models also tend to over-predict government-leaning stances, often misinterpreting ambiguous narratives. BanglaBias and its associated diagnostics provide a foundation for advancing stance detection in Bangla media research and offer insights for improving LLM performance in low-resource languages. [^1] https://anonymous.4open.science/r/BanglaBias
@@ 1 Introduction
News media often reflect subtle political framing and bias, shaping public opinion and behavior (Fuhat and Wahab, 2024). While extensive research has addressed this problem in English by creating rich datasets for sentiment analysis Elbouanani et al. (2025); Abercrombie and Batista-Navarro (2020), stance detection Rostami et al. (2025); Khiabani and Zubiaga (2025), hate speech identification Davidson et al. (2017), and political bias classification Jones (2024), the same cannot be said for many other languages. As Fig.˜1 illustrates, Bangla news media research faces a critical gap: despite clear evidence of political framing in coverage Islam (2016), there exists a scarcity of annotated datasets specifically designed to detect political stance and government affiliation bias. Our evaluations in Fig.˜1 revealed that while massive or proprietary models achieve strong performance on critique stance detection, neutral stance remains challenging, and Nano models showing large error bars underscores unstable performance in low-resource scenarios.
This disparity is particularly concerning given the widespread nature of political bias in Bangla news. Existing studies (both qualitative and computational) provide clear evidence that Bangla news content exhibits partisan frames that favor ruling-party narratives (Islam, 2016; Al-Zaman and Noman, 2024). For instance, two news outlets covering the same political event, such as a government policy announcement, may frame it as either a progressive reform or an authoritarian overreach, despite reporting the same factual details Vallejo et al. (2023). Such framing differences are not merely stylistic; they represent systematic ideological bias – a partisan slant away from neutral reporting (McQuail and Deuze, 2020).
Computational studies have begun to tackle this with machine learning, e.g., Bangla BERT Bhattacharjee et al. (2021) for hyper-partisanship Mehadi Hasan et al. (2025) and innovative LLM credibility bias Prama and Islam (2025). However, these efforts are hampered by the lack of comprehensive and well-annotated datasets that can evaluate robust detection of political positions, specifically for the Bangladeshi socio-political landscape. Audience studies further highlight how perceived slant and censorship are reshaping Bangla news consumption Islam et al. (2025).
@@
  • We introduce the first-ever benchmark dataset for political stance detection in Bangla with comprehensive metadata and a leaderboard [^2] Leaderboard will be published upon acceptance.. The dataset provides contextual information to study how political narratives are constructed in Bangladeshi media and enables the development and evaluation of stance detection models in such a low-resource setting.
@@ 2 Relevant Work
With the growing body of research on framing in NLP (Card et al., 2015; Fan et al., 2019; Baly et al., 2020; Ziems and Yang, 2021), and LLM prompting strategies (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023) to detect stances, recent studies have explored media bias through multiple computational approaches, as we discuss in the following sections.
@@ 2.1 Framing & Political Bias in News Media
Entman (2007) defines framing as the process of shaping narratives to promote specific interpretations, which then primes audiences to think in particular ways. Computational NLP work has built on these ideas by annotating and classifying news frames and biases (Card et al., 2015; Guida et al., 2025). Early efforts include the Media Frames Corpus (Card et al., 2015) and its multilingual extension (Piskorski et al., 2023), while others focus on ideological bias, such as BASIL (Fan et al., 2019), AllSides (Baly et al., 2020), and smaller resources on regional or issue-specific perspectives (Lin et al., 2006; Chen et al., 2018; Ziems and Yang, 2021). Such datasets do not capture the socio-political nuances of Bangladeshi cultural context, and there exists no such Bengali counterpart yet.
In practice, most NLP work treats media bias as a classification problem: one common approach is single-label classification of an article’s ideology. For instance, Recasens et al. (2013) derived word-level ideological features; Spinde et al. (2021) identified biased words via embedding distances between left- and right-leaning. Transformer-based models (e.g., BERT variants) now dominate framing tasks (Liu et al., 2019; Akyürek et al., 2020; Piskorski et al., 2023). Mehadi Hasan et al. (2025) presented a Bangla BERT model fine-tuned to identify hyperpartisan (extremely biased) news articles, achieving 95.7% accuracy, using a semi-supervised approach. The authors emphasized Bangla’s “low-resource” status – few annotated corpora, lexicons or pretrained models, and therefore relied on clever workarounds, i.e., machine translation (MT) for data augmentation. While MT is a practical approach, it fails to identify the socio-political facts that are specific to the region.
@@ 2.2 LLMs for Political & Subjective Analysis
Large pretrained language models have become popular for political stance and bias tasks via prompt or instruction-based methods. Researchers have evaluated models like GPT-3/GPT-4 and open alternatives (Falcon, LLaMA, Mistral, etc.) on stance detection and bias classification (Faulborn et al., 2025; Ng et al., 2025). For example, Ng et al. (2025) tested GPT-3.5 and seven open-source LLMs on three stance datasets (SemEval-2016 tweets (Mohammad et al., 2016), Elections-2016 (Sobhani et al., 2017) tweets, and the BASIL (Fan et al., 2019) news articles); they tried different prompting schemes (task description, context, chain-of-thought) and found reasoning based prompting scheme demonstrating the best performance in stance classification.
Sucu et al. (2025) showed that adding user-context information to prompts can boost performance: in a political forum stance task, contextual prompting improved accuracy by 17.5% to 38.5% over baseline. Lihammer (2023) showed GPT-3.5 and GPT-4 being able to reproduce political viewpoints. When tested in Bangladeshi context, most LLMs favored the left-leaning sources, giving higher credibility ratings (Prama and Islam, 2025). Instruction-tuned LMs can perform stance/bias tasks in a zero-shot or few-shot manner, but their outputs must be carefully validated as they may favor certain viewpoints and result in AI-induced bias in evaluating news sources, more significantly in low-resource languages like Bangla (Ng et al., 2025; Prama and Islam, 2025).
@@ 3 Creating BanglaBias
Prior research on Bangla social discourse classification (Haider et al., 2024) highlighted the challenges of transliteration noise, subtle category distinctions, and embedding reliability. Beyond these technical difficulties, a major bottleneck remains – the scarcity of high-quality, publicly available, politically nuanced Bangla datasets, which limits the development and evaluation of models for political stance detection and bias analysis. While translation or synthetic dataset generation has shown promise in domains like mathematics Toshniwal et al. (2024), instruction following Kim and Park (2024), or general language tasks Liu et al. (2024), these approaches are often ineffective for political stance detection and media bias analysis in Bangla as socio-political contexts are deeply tied to cultural, historical, and linguistic nuances that are lost or distorted in translation. Building on the research gap, we developed a systematic pipeline to create the first annotated Bangla political stance detection benchmark dataset.
@@
Annotator 1 and Annotator 2 achieved 73.5% agreement (Cohen’s $\kappa$ = 0.574), indicating moderate inter-rater reliability. Most of the disagreements arose from neutral vs government critique and some neutral vs government leaning confusions. Disagreements were resolved by a third annotator (adjudicator) after mutual discussion and reanalysis of the political affiliations of the respective news media.
@@ 4.1 Evaluation Metrics
We conduct evaluations in a reasoning-based detection setting. Following Ng et al. (2025), we prompted models to classify the article and provide a brief reasoning. This framework supports additional qualitative analyses to reveal performance differences across model sizes:
@@ 4.2 Models
  • Massive: Llama3.3-70B, GPT-OSS-120B, Qwen3-235B-I, Qwen3-235B-T, GLM-4.5, DeepSeek-V3.1, DeepSeek-R1, Grok-4.
  • Proprietary: Claude-Sonnet-4, Gemini-2.5-Pro, GPT-4.1-Mini, GPT-5-Mini.
Models are prompted to generate both a stance label (Decision: D) and reasoning $(R\toD)$. Full model ids can be found in App.˜E. The evaluated models include families such as Qwen Bai et al. (2023), LLaMA Touvron et al. (2023), GLM Zeng et al. (2025), DeepSeek Bi et al. (2024), GPT Achiam et al. (2023), Claude Caruccio et al. (2024), and Gemini Ng et al. (2025).
@@ 5.1 Performance Metrics Analysis
Although large-scale models (Massive and Proprietary) demonstrate clear dominance in classification tasks, standard models like Qwen3-32B and even smaller TigerLLM Raihan and Zampieri (2025) show competitive results in detecting stances, particularly government critique and neutral. This efficiency suggests that smaller models can effectively handle key linguistic features, such as critical keywords or neutral tone indicators, without requiring the computational resources of their larger counterparts. The strong performance of TigerLLM, in particular, highlights the potential of small domain-optimized models tailored for Bangla text, which benefit from focused training on regional linguistic patterns. In low-resource settings, where computational constraints are a significant concern, such resource-efficient models can be further optimized to perform critical political stance detection tasks.
@@ 5.2 Bias Tendency
Figure 4: Aggregated Confusion Heatmap over five categories of models: Nano (3 models), Compact (7 models), Standard (6 models), Massive (8 models) and Proprietary (4 models).
@@
The confusion heatmaps in Fig.˜4 reveal that smaller model categories like Nano and Compact frequently confuse neutral labels with government-leaning or critique, as seen in Nano’s 140 neutral instances misclassified as government-leaning and Compact’s 189 neutrals mistaken for government-critique. This indicates a bias toward polarized predictions due to limited capacity. As model scale increases, such as in Standard and Massive categories, mispredictions decrease, with Massive models showing balanced performance but still confusing neutral with government-critique. Proprietary models minimize cross-polarity errors (only 18 government-critique predicted as government leaning), yet persist in confusing neutral with polar stances (44 and 58 instances).
@@ 5.3 Error Analysis
We hypothesize that as the article praises such persona, it must mean a favoring stance. Predicting such article stance is much harder as it requires world knowledge and context outside the immediate text Burnham (2025).
@@ Appendix E Model ID Mapping
Grok-4 x-ai_grok-4-fast Massive

### v2

@@ Abstract
Detecting media bias is crucial, specifically in the South Asian region. Despite this, annotated datasets and computational studies for Bangla political bias research remain scarce. Crucially because, political stance detection in Bangla news requires understanding of linguistic cues, cultural context, subtle biases, rhetorical strategies, code-switching, implicit sentiment, and socio-political background. To address this, we introduce BanglaBias, the first benchmark dataset of 200 politically significant and highly debated Bangla news articles, labeled for government-leaning, government-critique, and neutral stances, alongside diagnostic analyses for evaluating large language models (LLMs). Our comprehensive evaluation of 28 proprietary and open-source LLMs shows strong performance in detecting government-critique content (F1 up to 0.83) but substantial difficulty with neutral articles (F1 as low as 0.00). Models also tend to over-predict government-leaning stances, often misinterpreting ambiguous narratives. BanglaBias and its associated diagnostics provide a foundation for advancing stance detection in Bangla media research and offer insights for improving LLM performance in low-resource languages. [^1] https://nusrat-lia.github.io/BanglaBias/
@@ 1 Introduction
News media often reflect subtle political framing and bias, shaping public opinion and behavior (fuhat2024mapping). While extensive research has addressed this problem in English by creating rich datasets for sentiment analysis elbouanani2025analyzing; abercrombie2020parlvote, stance detection rostami2025politisky24; khiabani2025cross, hate speech identification hateoffensive, and political bias classification cajcodes_political_bias, the same cannot be said for many other languages. As Fig.˜1 illustrates, Bangla news media research faces a critical gap: despite clear evidence of political framing in coverage islam2016coverage, there exists a scarcity of annotated datasets specifically designed to detect political stance and government affiliation bias. Our evaluations in Fig.˜1 revealed that while massive or proprietary models achieve strong performance on critique stance detection, neutral stance remains challenging, and Nano models showing large error bars underscores unstable performance in low-resource scenarios.
This disparity is particularly concerning given the widespread nature of political bias in Bangla news. Existing studies (both qualitative and computational) provide clear evidence that Bangla news content exhibits partisan frames that favor ruling-party narratives (islam2016coverage; al2024investigating). For instance, two news outlets covering the same political event, such as a government policy announcement, may frame it as either a progressive reform or an authoritarian overreach, despite reporting the same factual details vallejo2023connecting. Such framing differences are not merely stylistic; they represent systematic ideological bias – a partisan slant away from neutral reporting (mcquail2020mcquail).
Computational studies have begun to tackle this with machine learning, e.g., Bangla BERT bhattacharjee2021banglabert for hyper-partisanship mehadi2025bangla and innovative LLM credibility bias prama2025evaluating. However, these efforts are hampered by the lack of comprehensive and well-annotated datasets that can evaluate robust detection of political positions, specifically for the Bangladeshi socio-political landscape. Audience studies further highlight how perceived slant and censorship are reshaping Bangla news consumption islam2025conventional.
@@
  • We introduce the first-ever benchmark dataset for political stance detection in Bangla with comprehensive metadata and a leaderboard. The dataset provides contextual information to study how political narratives are constructed in Bangladeshi media and enables the development and evaluation of stance detection models in such a low-resource setting.
@@ 2 Relevant Work
With the growing body of research on framing in NLP (card2015media; fan2019plain; baly2020we; ziems2021protect), and LLM prompting strategies (brown2020language; achiam2023gpt; touvron2023llama) to detect stances, recent studies have explored media bias through multiple computational approaches, as we discuss in the following sections.
@@ 2.1 Framing & Political Bias in News Media
entman2007framing defines framing as the process of shaping narratives to promote specific interpretations, which then primes audiences to think in particular ways. Computational NLP work has built on these ideas by annotating and classifying news frames and biases (card2015media; guida2025retain). Early efforts include the Media Frames Corpus (card2015media) and its multilingual extension (piskorski2023semeval), while others focus on ideological bias, such as BASIL (fan2019plain), AllSides (baly2020we), and smaller resources on regional or issue-specific perspectives (lin2006side; chen2018learning; ziems2021protect). Such datasets do not capture the socio-political nuances of Bangladeshi cultural context, and there exists no such Bengali counterpart yet.
In practice, most NLP work treats media bias as a classification problem: one common approach is single-label classification of an article’s ideology. For instance, recasens2013linguistic derived word-level ideological features; spinde2021identification identified biased words via embedding distances between left- and right-leaning. Transformer-based models (e.g., BERT variants) now dominate framing tasks (liu2019detecting; akyurek2020multi; piskorski2023semeval). mehadi2025bangla presented a Bangla BERT model fine-tuned to identify hyperpartisan (extremely biased) news articles, achieving 95.7% accuracy, using a semi-supervised approach. The authors emphasized Bangla’s “low-resource” status – few annotated corpora, lexicons or pretrained models, and therefore relied on clever workarounds, i.e., machine translation (MT) for data augmentation. While MT is a practical approach, it fails to identify the socio-political facts that are specific to the region.
@@ 2.2 LLMs for Political & Subjective Analysis
Large pretrained language models have become popular for political stance and bias tasks via prompt or instruction-based methods. Researchers have evaluated models like GPT-3/GPT-4 and open alternatives (Falcon, LLaMA, Mistral, etc.) on stance detection and bias classification (faulborn2025only; ng2025examining). For example, ng2025examining tested GPT-3.5 and seven open-source LLMs on three stance datasets (SemEval-2016 tweets (mohammad2016semeval), Elections-2016 (sobhani2017dataset) tweets, and the BASIL (fan2019plain) news articles); they tried different prompting schemes (task description, context, chain-of-thought) and found reasoning based prompting scheme demonstrating the best performance in stance classification.
sucu2025exploiting showed that adding user-context information to prompts can boost performance: in a political forum stance task, contextual prompting improved accuracy by 17.5% to 38.5% over baseline. lihammer2023semantic showed GPT-3.5 and GPT-4 being able to reproduce political viewpoints. When tested in Bangladeshi context, most LLMs favored the left-leaning sources, giving higher credibility ratings (prama2025evaluating). Instruction-tuned LMs can perform stance/bias tasks in a zero-shot or few-shot manner, but their outputs must be carefully validated as they may favor certain viewpoints and result in AI-induced bias in evaluating news sources, more significantly in low-resource languages like Bangla (ng2025examining; prama2025evaluating).
@@ 3 Creating BanglaBias
Prior research on Bangla social discourse classification (haider2024banth) highlighted the challenges of transliteration noise, subtle category distinctions, and embedding reliability. Beyond these technical difficulties, a major bottleneck remains – the scarcity of high-quality, publicly available, politically nuanced Bangla datasets, which limits the development and evaluation of models for political stance detection and bias analysis. While translation or synthetic dataset generation has shown promise in domains like mathematics toshniwal2024openmathinstruct, instruction following kim2024instatrans, or general language tasks liu2024translation, these approaches are often ineffective for political stance detection and media bias analysis in Bangla as socio-political contexts are deeply tied to cultural, historical, and linguistic nuances that are lost or distorted in translation. Building on the research gap, we developed a systematic pipeline to create the first annotated Bangla political stance detection benchmark dataset.
@@
Annotator 1 and Annotator 2 achieved 73.5% agreement (Cohen’s $\kappa$ = 0.574), indicating moderate inter-rater reliability. Most of the disagreements arose from neutral vs government critique and some neutral vs government leaning confusions. Disagreements were resolved by a third annotator (adjudicator) after mutual discussion and reanalysis of the political affiliations of the respective news media. Annotators primarily adhered to the stance detection decision flow outlined in App.˜H.
@@ 4.1 Evaluation Metrics
We conduct evaluations in a reasoning-based detection setting. Following ng2025examining, we prompted models to classify the article and provide a brief reasoning. This framework supports additional qualitative analyses to reveal performance differences across model sizes:
@@ 4.2 Models
  • Massive: Llama3.3-70B, GPT-OSS-120B, Qwen3-235B-I, Qwen3-235B-T, GLM-4.5, DeepSeek-V3.1, DeepSeek-R1.
  • Proprietary: Claude-Sonnet-4, Grok-4, Gemini-2.5-Pro, GPT-4.1-Mini, GPT-5-Mini.
Models are prompted to generate both a stance label (Decision: D) and reasoning $(R\toD)$. Full model ids can be found in App.˜E. The evaluated models include families such as Qwen bai2023qwen, LLaMA touvron2023llama, GLM zeng2025glm, DeepSeek bi2024deepseek, GPT achiam2023gpt, Claude caruccio2024claude, and Gemini ng2025examining.
@@ 5.1 Performance Metrics Analysis
Although large-scale models (Massive and Proprietary) demonstrate clear dominance in classification tasks, standard models like Qwen3-32B and even smaller TigerLLM raihan2025tigerllm show competitive results in detecting stances, particularly government critique and neutral. This efficiency suggests that smaller models can effectively handle key linguistic features, such as critical keywords or neutral tone indicators, without requiring the computational resources of their larger counterparts. The strong performance of TigerLLM, in particular, highlights the potential of small domain-optimized models tailored for Bangla text, which benefit from focused training on regional linguistic patterns. In low-resource settings, where computational constraints are a significant concern, such resource-efficient models can be further optimized to perform critical stance detection tasks.
@@ 5.2 Bias Tendency
Figure 4: Aggregated Confusion Heatmap over five categories of models: Nano (3 models), Compact (7 models), Standard (6 models), Massive (7 models) and Proprietary (5 models).
@@
The confusion heatmaps in Fig.˜4 reveal that smaller model categories like Nano and Compact frequently confuse neutral labels with government-leaning or critique, as seen in Nano’s 140 neutral instances misclassified as government-leaning and Compact’s 189 neutrals mistaken for government-critique. This indicates a bias toward polarized predictions due to limited capacity. As model scale increases, such as in Standard and Massive categories, mispredictions decrease, with Massive models showing balanced performance but still confusing neutral with government-critique. Proprietary models minimize cross-polarity errors (only 25 government-critique predicted as government leaning), yet persist in confusing neutral with polar stances (55 and 65 instances).
5.3 Prompting Strategy Comparison:
To further investigate the impact of prompting on stance detection, we compare zero-shot (no example) and few-shot (multiple examples) prompting strategies across selected models, focusing on those with varying sizes (e.g., 1.7B to 70B parameters).
In subjective tasks like political bias classification in Bangla news articles, few-shot prompting can underperform compared to zero-shot, particularly in larger models (see App.˜I), due to risks of overfitting to example patterns, such as over-emphasizing neutrality if examples are imbalanced or create conservative heuristics that dismiss implicit biases as “factual.” This leads to systematic errors like defaulting to “Neutral” labels, as seen in aggregate metrics where few-shot Weighted F1 scores lag (e.g., 0.57–0.67 vs. zero-shot’s 0.68–0.73). However, few-shot may still edge out in smaller models (e.g., 1.7B) needing more examples for calibration. This highlights that prompting-strategy efficacy depends on model scale, multiple patterns of framing and dataset nuances like implicit political tones.
@@ 5.4 Error Analysis
We hypothesize that as the article praises such persona, it must mean a favoring stance. Predicting such article stance is much harder as it requires world knowledge and context outside the immediate text burnham2025stance.
@@ Appendix E Model ID Mapping
Grok-4 x-ai_grok-4-fast Proprietary
@@ Appendix F Human Annotator Details
Appendix G Evaluation Prompt
Zero-shot evaluation prompt You are a political bias classifier. Task: Read the following article and return ONLY a valid JSON object with two fields: {{ "label": one of ["Govt leaning", "Neutral", "Govt critique"], "reason": "1–3 sentences explanation" }} Rules: - Output ONLY JSON. - Do not include markdown, code fences, or explanations. Article: {{article_text}}
Few-shot evaluation prompt You are a political bias classifier. Task: Read the following article and return only a valid JSON object with two fields: {{{{ "label": one of ["Govt leaning", "Neutral", "Govt critique"], "reason": "1–3 sentences explanation" }}}} Rules: Output ONLY JSON (no markdown, no extra text). Base your decision strictly on tone, stance, and framing not just factual accuracy. Examples: Example 1: Article: ... ... Output: {{"label":"Neutral","reason":"..."}} Example 2: Article: ... ... Output: {{"label":"Govt leaning","reason":"..."}} Example 3: Article: ... ... Output: {{"label":"Govt critique","reason":"..."}} Now classify the following article: {article_text}
Appendix H Annotation Decision Flowchart
A high-level annotation decision flowchart is provided in the Fig.˜6.
Figure 6: Annotation Decision Flowchart for Political Stance Classification in the Bangladeshi Context. The decision flowchart operationalizes stance annotation within a politically polarized media system, where neutrality is a negotiated discursive position rather than a natural default.
Appendix I Zero-shot vs Few-shot
Zero-shot vs few-shot prompting is provided in Table˜6.
Model Govt. Critique Neutral Govt. Leaning Weighted Avg
P R F1 P R F1 P R F1 P R F1
Qwen3-1.7B (Few-Shot) 0.66 0.78 0.71 0.56 0.60 0.58 0.55 0.18 0.27 0.60 0.62 0.59
Qwen3-1.7B (Zero-Shot) 0.71 0.69 0.70 0.54 0.21 0.30 0.28 0.67 0.39 0.58 0.52 0.51
TigerLLM-9B (Few-Shot) 0.86 0.58 0.69 0.51 0.82 0.63 0.57 0.36 0.44 0.69 0.63 0.63
TigerLLM-9B (Zero-Shot) 0.74 0.78 0.76 0.62 0.65 0.64 0.62 0.45 0.53 0.68 0.68 0.68
Qwen3-32B (Few-Shot) 0.84 0.45 0.59 0.45 0.76 0.57 0.54 0.45 0.49 0.65 0.57 0.57
Qwen3-32B (Zero-Shot) 0.82 0.68 0.75 0.55 0.72 0.63 0.70 0.58 0.63 0.71 0.68 0.68
Llama3.3-70B (Few-Shot) 0.86 0.68 0.76 0.59 0.57 0.58 0.49 0.82 0.61 0.70 0.67 0.67
Llama3.3-70B (Zero-Shot) 0.77 0.91 0.83 0.80 0.50 0.62 0.57 0.76 0.65 0.75 0.73 0.73
Table 6: Few-Shot vs Zero-Shot Performance Comparison